OpenStudy (anonymous):

Solve the equation. Check both solutions and only write the real solution(s).

OpenStudy (anonymous):

\[\sqrt{2X+3}=X\]

OpenStudy (anonymous):

2x+3 = x^2 x^2-2x-3=0 (x+1)(x-3) x=-1 & x = 3